# WaterVG: Waterway Visual Grounding based on Text-Guided Vision and mmWave Radar

### v1

@@ 5.2. Quantitative Results
Table 1 presents the overall performances of our proposed Potamoi and other competitors on the benchmarks of VVG and FVG. Firstly, overall, we find that all models perform better on FVG than on VVG. This suggests that relying solely on visual features may not accurately localize specific described targets, particularly those incorporating distance or motion features in textual descriptions. Secondly, our proposed Potamoi achieves competitive performance compared to other visual grounding models, whether based on transformer architecture (TransVG, SeqTR, RefTrans) or one-stage (MCN, SimREC, LiteREC-S2) or two-stage paradigms (MattNet). Specifically, for the REC task of VVG and FVG, our Potamoi achieves state-of-the-art performances overall, which in a race with RefTrans (Li and Sigal, 2021) and SeqTR (Zhu et al., 2022). Moreover, for RES task, due to the lightweight segmentation head (Peng et al., 2022) utilized in our Potamoi, it generally ranks second or third among all models in Table 1, trailing behind the dedicated RES model VLT (Ding et al., 2022) based on transformer architecture with a larger parameter count. Nevertheless, from the perspective of multi-task models, our Potamoi overperforms MCN, SeqTR and is evenly contested to RefTrans but a little better.
5.3. Qualitative Results
Figure 8 visualises REC and RES results of our proposed Potamoi and MCN (Luo et al., 2020) on the representative samples of WaterVG, including (a) piecewise multi-objective prompt, (b) distinguish one of multiple objects based on the description of the target distance, (c) multi-object localization based on fuzzy text prompts and (d) count-based fuzzy hints. It can be observed that our proposed Potamoi achieves more prominent performance compared to MCN, obtaining lower target miss rates, false positive rates, and masks that better fit the shapes and contours of the targets.
5.4. Ablation Experiments
@@
MHCA (Dao et al., 2022) 66.048K 293.27M 88.00 71.8 70.23
0.263M 257.23M 64.06
1.050M 256.41M 60.00
MHLCA (Choromanski et al., 2020) 0.252M 1.691G 124.02 70.2 72.01
0.524M 865.08M 94.27
1.049M 471.86M 45.74
@@
MHSCA 0.768K 8.14M 108.53 71.2 71.04
1.536K 4.07M 66.14
3.072K 2.04M 44.29
@@
Cross attention comparison. In Table 2, we conduct a comprehensive comparison between our MHSCA and two representative multi-head cross-attention mechanisms, MHCA (Dao et al., 2022) and MHLCA (Choromanski et al., 2020), which are transplanted to our Potamoi, encompassing comparisons across different channel dimensions and spatial sizes of visual feature maps. Moreover, we adjust the agent number $n$ in MHSCA as 49 and 144 for comparison, respectively.
@@
Firstly, looking at the performance on both REC and RES tasks, our proposed MHSCA ($n$=144) achieves competitive results. It outperforms MHLCA slightly in the REC task and lags behind MHCA, while in the RES task, it surpasses MHCA but still trails MHLCA. However, it is noteworthy that for visual feature maps with increasing spatial sizes, the FLOPs of our MHSCA never exceed 10M and demonstrate smaller growth compared to MHCA and MHLCA. Regarding expanding channels, our MHSCA maintains a significantly lower parameter count compared to the other two fusion methods. Importantly, we also compare the power consumption of three fusion methods, where our MHSCA with $n$=49 achieves the lowest value for three-dimensional feature maps. MHSCA with $n$=144 obtains competitive performance with MHCA based on flash attention (Dao et al., 2022). Therefore, our MHSCA achieves a well-balanced multi-modal fusion approach between accuracy, complexity and power consumption, serving as a lightweight cross-attention mechanism.
@@
Comparison of other settings. As Table 3 shows, when adopting multi-task learning, the performance of Potamoi is apparently better than single-task learning, which indicates multi-task learning can jointly promote the performances of individual tasks. Figure 9 visualizes the last stage of P4 in the Lite-PAN before REC and RES heads, we find that Potamoi successfully focuses on the spatial locations of one or more targets referred to in textual prompts, indicating that REC and RES are optimized in the right direction. Table 4 presents the ablation experiments of other settings in fusion-based Potamoi, mainly including the settings of input data (RVP maps) and neural networks. We find that in the radar features of targets, target distance and velocity are relatively more vital for visual grounding. Additionally, due to some samples having text prompt token lengths exceeding 30, and certain target-descriptive terms occurring towards the end of text, reducing token length resulted in information loss, leading to a decrease in prediction accuracy. Moreover, Adaptive Radar Weighting (ARW), as well as SPPM, play significant roles on feature representation.
Table 3. Ablation experiments on multi-task learning.
[…]
70.9 44.0 52.6 70.68
70.1 44.8 53.0 70.81
@@
No SPPM - - - 69.52
No Query Pos 68.9 42.7 52.0 69.97
Figure 9. Last-stage heatmaps of PAN before prediction heads.
@@ 6. Conclusion
In this paper, we propose a new visual grounding dataset, WaterVG, for intelligent perception in waterway environments, which includes textual prompts, visual, and millimeter-wave radar data. Additionally, each sample’s text prompt contains references to multiple targets. Based on this, we design an efficient multitask visual grounding model, Potamoi. It is capable of accurately performing REC and RES tasks simultaneously based on textual prompts containing target appearance, position, distance features, and more. This capability relies on our proposed lightweight and highly effective Multi-Head Slim Cross Attention (MHSCA) for effective fusion of textual, visual, and radar features. Both Potamoi and MHSCA demonstrate competitive performance compared to counterparts.

### v2

@@ 5.2. Quantitative Results
Table 1 presents the overall performances of our proposed Potamoi and other competitors on the benchmarks of VVG and FVG. Firstly, overall, we find that all models perform better on FVG than on VVG. This suggests that relying solely on visual features may not accurately localize specific described targets, particularly those incorporating distance or motion features in textual descriptions. Secondly, our proposed Potamoi achieves competitive performance compared to other visual grounding models, whether based on transformer architecture (TransVG, SeqTR, RefTrans) or one-stage (MCN, SimREC, LiteREC-S2) or two-stage paradigms (MattNet). Specifically, for REC task of VVG and FVG, our Potamoi achieves state-of-the-art performances overall, which in a race with RefTrans (Li and Sigal, 2021) and SeqTR (Zhu et al., 2022). Moreover, for RES task, due to the lite segmentation head (Peng et al., 2022) in Potamoi for power saving, it generally ranks second or third among all in Table 1, trailing behind the dedicated RES model VLT (Ding et al., 2022) based on transformer architecture with a larger parameter count. Nevertheless, among multi-task models, our Potamoi overperforms MCN, SeqTR and is evenly contested to RefTrans. Table 2 presents the performance of Potamoi on test set under different scenarios and sentence query patterns as Figure 3 presents. Importantly, our Potamoi achieves the lowest power consumption for both VVG or FVG, which is much lower than both single-task and multi-task models, implying Potamoi can help extend the endurance time of USV operations. Lastly, Table 4 and 5 show that our low-power Potamoi obtains competitive performances compared with other state-of-the-art models on RefCOCOs. Further, our MHSCA can be nicely adapted to CNN-based MCN and transformer-based SeqTR to improve performance.
@@
MHCA (Dao et al., 2022) 66.048K 293.27M 71.8 45.0 52.9 70.23
0.263M 257.23M
1.050M 256.41M
MHLCA (Choromanski et al., 2020) 0.252M 1.691G 70.2 44.7 53.2 72.01
0.524M 865.08M
1.049M 471.86M
@@
MHSCA 0.768K 8.14M 71.7 45.1 53.1 71.09
1.536K 4.07M
3.072K 2.04M
Length of Sensor Agent $n$ = 256 (16 $\times$ 16)
MHSCA 0.768K 8.24M 72.3 45.2 54.0 71.95
1.536K 3.98M
3.072K 2.26M
Figure 9. Prediction samples of Potamoi on four primary query types of textual prompts.
5.3. Qualitative Results
Figure 8 visualizes predictions of our Potamoi and MCN (Luo et al., 2020) on the representative samples of WaterVG, including (a) distinguish partial targets based on the query of target distance, (b) multi-target localization based on fuzzy text prompts and (c) count-based fuzzy hints. Apparently, our Potamoi behaves more prominent than MCN, obtaining lower target miss rates, false positive rates, and masks that better fit shapes and contours of the targets. Moreover, Figure 9 presents prediction samples upon four primary query types in WaterVG, where our Potamoi exhibits nice adaptation.
5.4. Ablation Experiments
@@
Cross attention comparison. In Table 3, we conduct a comprehensive comparison between our MHSCA and two representative multi-head cross-attention mechanisms, MHCA (Dao et al., 2022) and MHLCA (Choromanski et al., 2020), which are transplanted to our Potamoi, encompassing comparisons across different channel dimensions and spatial sizes of visual feature maps in PAN. Moreover, we adjust the agent number $n$ in MHSCA as 49, 144 and 256 for comparison, respectively.
@@
Firstly, looking at the performance on both REC and RES tasks, our proposed MHSCA ($n$=256) achieves the overall best results. MHSCA ($n$=144) also obtains competitive performances. It is noteworthy that for visual feature maps with increasing spatial sizes, the FLOPs of our MHSCA never exceed 10M and demonstrate smaller growth compared to flash attention-based (Dao et al., 2022) MHCA and MHLCA. Regarding expanding channels, our MHSCA maintains a significantly lower parameter count compared to the other two fusion methods. Further, Figure 10 visualizes the last stage of PAN after fusing multi-scale attention maps under settings of MHSCA and MHCA, MHSCA correctly capture targets referred by prompt, also advanced in small target localization. Therefore, MHSCA achieves a well-balanced result between accuracy and complexity, serving as a lite and scalable cross-attention method for multi-scale maps.
Figure 10. Last-stage heatmaps of PAN before prediction heads.
Table 4. Comparison of REC on RefCOCOs datasets.
Models RefCOCO RefCOCO+ RefCOCOg
testA testB testA testB test-u
MAttNet 80.43 69.28 70.26 56.00 67.27
TransVG 82.72 78.35 70.70 56.94 67.73
RefTrans 85.53 76.31 75.58 61.91 69.10
MCN 82.29 74.98 72.86 57.31 66.01
MCN-MHSCA 83.06$\uparrow$ 75.22$\uparrow$ 73.54$\uparrow$ 58.67$\uparrow$ 66.92$\uparrow$
SeqTR 85.00 76.08 75.37 58.78 71.58
SeqTR-MHSCA 85.36$\uparrow$ 76.72$\uparrow$ 75.85$\uparrow$ 59.10$\uparrow$ 72.13$\uparrow$
Potamoi 84.98 78.10 75.90 58.55 70.55
Table 5. Comparison of RES on RefCOCOs datasets.
Models RefCOCO RefCOCO+ RefCOCOg
testA testB testA testB test-u
MAttNet 62.37 51.70 52.39 40.08 48.61
 VLT 68.29 62.73 59.20 49.36 56.65
MCN 64.20 59.71 54.99 44.69 49.40
MCN-MHSCA 64.51$\uparrow$ 59.94$\uparrow$ 55.30$\uparrow$ 45.32$\uparrow$ 49.76$\uparrow$
SeqTR 69.79 64.12 58.93 48.19 55.64
SeqTR-MHSCA 70.10$\uparrow$ 64.32$\uparrow$ 59.09$\uparrow$ 48.33$\uparrow$ 55.90$\uparrow$
Potamoi 68.71 63.02 58.70 48.69 54.12
@@
Comparison of other settings. Table 6 presents ablation experiments of other settings in fusion-based Potamoi, including settings of input data, network and training. For radar features, target range and velocity are more vital for visual grounding. Besides, due to some samples containing prompt token lengths exceeding 30, and certain target-descriptive terms occurring towards the end of text, reducing token length resulted in information loss, leading to a decrease in accuracy. Moreover, when removing ARW and only with element-wise addition for vision-radar fusion, an apparent drop appears. When adopting multi-task learning, the performance of Potamoi is better than single-task learning, which indicates multi-task learning can promote performances of both tasks.
@@
Training
Single-task 70.9 44.0 52.6 70.68
@@ 6. Conclusion
We propose the first visual grounding dataset of USV-based waterway perception, WaterVG, which includes fine-grained textual prompts, visual, and millimeter-wave radar data. Further, each sample’s text prompt contains references to multiple targets. Based on this, we design a multi-task visual grounding model, Potamoi, which can accurately performing REC and RES tasks simultaneously with the lowest power consumption among all models. This capability relies on our proposed ARW and highly efficient Multi-Head Slim Cross Attention (MHSCA) for fusion of textual and sensor features with a considerably low cost. Both Potamoi and MHSCA demonstrate competitive performance compared to counterparts.